The P/E ratio on a stock market is 14. The underlying real earnings growth rate is 1.4%, at a constant long-run inflation rate of 2.7%. What is the firm's nominal long-run cost of capital?
Answer #1

Real rate = 1/(PE) + g = 1/14 + 0.014 = 0.0854 or 8.54%

Nominal rate = [( 1 + real rate ) x ( 1 + Inflation rate )] - 1

= [ 1.0854 x 1.027 ] - 1

= 0.1147 or 11.47%